With dsDNA (4.4 or 1.1 kb): good results almost every time, Good tether: stuck ratio, Good over all tether density (a couple / field of view)
Good monodisperse beads, Dilution technique, Sonication conditions
Good sample chambers, cleaning glass necessary, sonicate in alconox and water for 3min and rinsing numerous times. Blow dry.
Good blocker, prevents stuck beads, still permits tethering
Good antibody, diluted with PBS, tethers possible
DNA Constructs, Plasmid only construct, Success via gel observation, success via tethering?, not unzipped yet, shotgun clones construct, Success via gel observation?, don't think so, without concatamers, Not necessary for biophys poster. Necessary for CAREER and other grants
Optical Tweezers, Hardware, Necessary tasks, AOM, Pranav has aligned AOM crystal and the crystal seems to work very well, The driver works well in CW mode, but not in "normal" mode., In communication w/ tech support, Tech support asked to send it back, Steve wants to ask that they send us a replacement first, so OT can be aligned in the meantime., Engineer called me, but I had my meeting with DARPA, so couldn’t continue the talk. Basically, we’re at a bit of an impasse. He says it works just fine and he can control it in “normal mode” and get 2 W power. I’ll talk with him more tomorrow morning, and if can’t make progress, we’ll need to do some thinking., Driver has been updated to work with 5V input., Just need to test now to see if get high efficiency diffraction in "normal" mode and then we're done., Beam Diameter, We never ascertained the optimal beam diameter (beam expander setting), and even then, it may be different with the new beam path, Can test via power spectrum at different beam expander settings, Drift problem, Seems like some components (perhaps the periscope) are drifting, Calibration, Read Richard Yeh's thesis for good background on this process, Detector sensitivity, mV / nm, Stuck bead sweeps, Very sensitive to bead height, Can be studied by adjusting laser trap position relative to focal plane, via the z-position of 1st telescope lens, For "quick" calibration, we need to know sensitivity where the bead will be (pretty much trap center, I think), Detector stiffness, pN / nm / mW, Free bead power spectrum, Have software for this (integrated with Versatile Feedback OT control program, Free bead, piezo triangle wave, Good for double-check, Laser problem, Call Company to see if can be retuned on-site by local service, See if cold aluminum block helps, Things to probably purchase, Better Translaters for the steering telescope, Better periscope mounts? (to prevent drift), Better condenser?, Software, OT Control code, Probably need some cleaning up and at the very least a good once over. It has been a long time since Larry looked at it., Data analysis code (converting raw to force versus unzipping index), In pretty good shape, but calibration will change